OpenStudy (anonymous):

iranmeah91 F.dA (FLUX) explaination

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hey explain what E.dA is in physics if you can compared to flux in mathematics

OpenStudy (anonymous):

they are the same

OpenStudy (anonymous):

in the books its E.dA but theres a 1/4pi(something else)

OpenStudy (anonymous):

This is electric Field\[E=\frac{ q}{4 \pi \epsilon_0 r^2}\] Flux in sphere E* A= \[E=\frac{ q}{4 \pi \epsilon_0 r^2}* 4\pi r^2\] \[E*A=\frac{ q}{4 \pi \epsilon_0 r^2}* 4\pi r^2\] \[E*A=\frac{ q}{\epsilon_0}\]

OpenStudy (anonymous):

ok kinda clears it up, but solving them you would tackle them same way?

OpenStudy (anonymous):

yes, normally you use Gaussian Surface to avoid using Surface Integral

OpenStudy (anonymous):

cool thanks, and same with stokes?

OpenStudy (anonymous):

if it is not sphere or cylinder , you would have to use those stuff(Div, Stoke)

OpenStudy (anonymous):

That's why I asked that question to krebante

OpenStudy (anonymous):

yeah, i was just seeing who knows their stuff here, i knew how to do flux n curl and grad very well in MATH, but i have not taken physics yet, which will prob be a little difficult. I just wanted to see who i could go to for physics questions instead of just posting top everyone

OpenStudy (anonymous):

applying those can be tricky at times

OpenStudy (anonymous):

but i have not done any of those yet in a physics context

OpenStudy (anonymous):

yeah ill def ask you if i got questions, thanks.